Topic
V July 28, 2015 - Personalization Stand-Up
* Attendees: Anne, Rajkamal, Anjali, Rajesh, Jagadish, Andrew, Kavida, Brian
* Phone Attendees: Matt K., Shankar, Archana, Vaidhya
V Hot Topics (In addition to status below)
* Announcement: Stand up will be 30 minutes earlier on Wednesday
* MDM Deletes - when do they need to go in? Before 8/13? — Kavida, Shankar, and Jagadish agree — not critical before 8/13
* Have we reviewed the VPD change before the deploy today to make sure there are no “production mismatches”? Does a checklist for this exist? — Checklist exists, Rajesh & Anjali will run through it. Recently expanded to catch the problem we had after the initial Res/Shops deploy
* UAT testing pending - test data issue, what remains from teams — Andrew to get feedback from teams today
V Status
V Analysis
V Jeff
* Out sick today
V Kavida
* Got code together for party and demographics. Made some updates to hopefully exercise demographics incrementals. This will allow us to burn down some additional story points in this sprint.
* Checked on the fix for “date of birth” issue (between GOMaster and MDM) deployed to Sleepy by DI sustainment team. Some additional work still pending.
* One time data fix by managed services team also looking good.
* Monitoring daily MDM incrementals to make sure they’re working well.
V Shankar
* Looking at reservation “primary guest” — source data issue. Will follow up with Nag to see if there is any further info/clarification.
V Anne
* Waiting on decision on using XI as a data source for Fastpass.
* Andrew’s request — look at planning for next Sprint with Jarrad and Jagadish
V Modeling
V Ranjit
* Continued updates to the Reservation extract mapping
* Analysis on the Tickets for sourcing from XI schema (Based on the inputs provided by Brain Normant)
V ETL
V Anjali
V Email updates:
V Production:
* VPD- Production deployment task created for 07/28. All deployment artifacts placed. Need confirmation if we can deploy DVIC along with VPD.
* PER-2159 - Guest XREF Extract observation on Incremental data- Fix required. Deployed. Closed
* MDM Deletes Changes- Design change required- Development in progress
* Production Historical run in Doc- Forklifting of data- Testing using Dev and SIT - In Progress
* Bteq changes for all the Reservation and Shops Work in Progress.
* MDM - SHA Mappings- Fast load changes– Planned for deployment today 07/28
* Changing commit Interval from 10000 to 1000000 for all the Fast Load workflows – MDM IA Components-Planned for deployment today 07/28
V Rajesh
* Worked with Sean on initial data load. Fixed one issue overnight… should wrap up the testing today.
* Looking at tuning long-pole mappings for Reservations.
V QA
V Rajkamal
* In Doc: loaded everything that was published yesterday. Paused until the deployments planned for today (see above) are complete.
* DVIC - files created for historical/incremental, testing in progress, should finish today.
* Demographics in Doc - counts a little bit off, investigating this today.
* With Rajesh, going to reach out to MDM (Dinesh and Khizer to see about using file-based triggers between UC4 Prod/Doc)
V Brian
* On XI all day yesterday
V DBA
V John
* Restore/rebuild of Reservation into Doc went fine so far, in progress today. No ETA
V PM
V Jagadish
* Lining out plan/timeline of performance tuning changes.